OpenStudy (anonymous):

Solve 12x2 – 28x + 4 = 0.

OpenStudy (anonymous):

x = 1/6 (7-sqrt(37)) x = 1/6 (7+sqrt(37)) a = 12 b = -28 c = 4 Use the Quadratic Formula